# From Ferminet to PINN. Connections between neural network-based algorithms for high-dimensional Schrödinger Hamiltonian

### v1

@@ Abstract
In this note, we establish some connections between standard (data-driven) neural network-based solvers for PDE and eigenvalue problems developed on one side in the applied mathematics and engineering communities (e.g. Deep-Ritz and Physics Informed Neural Networks (PINN)), and on the other side in quantum chemistry (e.g. Variational Monte Carlo algorithms, Ferminet or Paulinet). In particular, we re-formulate a PINN algorithm as a fitting problem with data corresponding to the solution to a standard Diffusion Monte Carlo algorithm initialized thanks to a neural network-based Variational Monte Carlo (e.g. Ferminet). Connections at the level of the optimization algorithms are also established.
@@ 1 Introduction
In less than 10 years (since 2017), Physics Informed Neural Networks (PINN) has emerged as a major player in the field of computational methods for solving direct and inverse problems involving (high-dimensional) parameterized partial differential equations (PDEs) [21, 25, 18, 1]. The principle is to seek solutions to PDEs in the form of data-driven neural networks, optimized by minimizing PDE residuals. These techniques, along with others such as polynomial regression, Gaussian process, reduced order modeling, or Deep Ritz [23], have become highly popular in several engineering fields (e.g., mechanical engineering, fluid dynamics) as they allow to a certain extent, the tackling of problems that are inaccessible with traditional PDE solvers.
@@
Independently of the development of PINN-like methods by the mathematics, engineering, computer science, and statistics communities, significant practical progress has been achieved in quantum chemistry and quantum physics in the development of neural network-based Variational Monte Carlo (VMC) and Diffusion Monte Carlo (DMC) methods. This has enabled the calculation of the electronic structure of molecules larger than those usually considered with traditional computational methods. This approach corresponds mathematically to the approximation of the lowest eigenvalues (energies) of high-dimensional Schrödinger operators. Several corresponding libraries have been developed, such as Ferminet by DeepMind///Google [9, 20] or Paulinet ([14]) following et al. [20, 7, 8, 24].
To our knowledge, the two above formalisms have never been unified or even rigorously connected, despite the strong similarities. In this manuscript, we “simply” establish some connections potentially relevant for the development of a mathematical formalism unifying these modern techniques. More specifically, we will reformulate the so-called Neural-Variation Diffusion Monte Carlo (N-VDMC), corresponding to a standard DMC method where the trial wavefunction is computed through a VMC method (commonly called Deep-Ritz algorithm in the applied mathematics community [23]) as proposed in Ferminet. We will also show that at the level of the optimization algorithm (gradient descent-like algorithm) a N-VDMC is indeed “similar” to a PINN method applied to a diffusion equation (PINN-DE) derived from the time-dependent Schrödinger equation.
@@
In this manuscript, we do not discuss the (very complex) structure of the neural networks usually used in quantum chemistry which follows standard strategies based on one-electron orbitals and Slater’s determinants [22]. We do not detail this construction and refer to [20, 7, 8] for details. However, the properties and connections presented in this manuscript are largely independent of the structure of the neural networks; the latter will be simply treated here as a parameterized functions.
@@ 1.2 Basics on neural networks
where $\sigma$ is a so-called activation function. We refer to [15, 2] for details on neural networks. The use of neural network-based algorithms for PDE and eigenvalue problems largely rely on approximability theorems such as the celebrated Cybenko’s Universal Approximation theorem [11] and more recently [12, 13]. Let us repeat that here, we do not discuss the very ellaborated structure of neural networks used for approximating wavefunctions in high-dimensional electronic structure problems, and refer to [20, 7, 8] for details.
@@ 1.3 Deep-Ritz /// Ferminet
Ferminet [9, 20] has become a standard electronic structure calculation library, which is based on the minimization of the Rayleigh quotient
@@
hence defining an approximate groundstate. This approach is referred as Deep-Ritz in the applied mathematics community [23]. The specificity of Ferminet and [7, 8, 20] is the very complex structure of the neural network (anti-symmetric, etc). In the following, the use of neural network as ansatz for minimizing the Rayleigh quotient will be referred as Neural-Variational Monte Carlo (N-VMC).
@@ 1.4 Diffusion Monte Carlo and Neural Diffusion Monte Carlo
leading to $E_{G}=\langle H\phi_{G},\psi_{T}\rangle/\langle\phi_{G},\psi_{T}\rangle$, as long as $\psi_{T}$ has an overlap with $\phi_{G}$. From a practical point of view the above approach is discretized using random walks. A sequence of time-dependent random walks $\{R^{\alpha}(t)\}_{\alpha}$ from $t$ to $\mathbb{R}^{3N}$ allows for the construction of a function $\widetilde{f}$ approximating $f$ in the form (see [26])
@@
A sequence of discrete times $t_{0}=0<t_{1}<\cdots<t_{i}<t_{i+1}<\cdots$ and $dt=t_{i+1}-t_{i}$ for any $i\in\mathbb{N}$ is introduced. For $\{R_{0}^{\alpha},t_{0}\}_{\alpha}$ given, the following sequence is then constructed
$\widetilde{f}_{i}^{\alpha}=\widetilde{f}(R_{i}^{\alpha},t_{i}),$
such that
@@
Finally, for $i$ large enough and denoting by $\langle\cdot\rangle_{\alpha}$ the average over the walkers, we get the mixed estimator at any time $t_{i}$, $E^{i}_{\textrm{mix}}=\langle E_{i-1}(R_{i}^{\alpha})\widetilde{f}_{i}^{\alpha}\rangle_{\alpha}/\langle\widetilde{f}_{i}^{\alpha}\rangle_{\alpha}$.
@@ 1.5 Basics on PINN
Physics Informed Neural Networks (PINN) [21, 25, 18, 1] has become a standard methodology in engineering and applied mathematics for data-driven computation of partial differential equations. A PINN-like solution to a differential system ($\mathcal{S}$) is a (smooth and possibly data-driven) neural network optimized thanks to the minimization of the norm of the PDE residual of ($\mathcal{S}$) evaluated at some (randomly chosen) learning nodes. This approach has recently gained a huge interest in scientific machine learning thanks to their simplicity, their nice mathematical and computational properties coming in particular from automatic differentiation and efficient stochastic gradient descent methods [5].
We now discuss the principle of PINN algorithms for solving (2). In order to avoid a null solution (which can occur for large times), it is preferable to decompose the time-domain as follows: $\mathbb{R}_{+}=\cup_{i=1}^{\infty}[t_{i-1},t_{i})$, with $t_{0}=0<t_{1}<\cdots<t_{i}<t_{i+1}<\cdots$ and we denote $dt=t_{i+1}-t_{i}$ and $Q_{i}=\mathbb{R}^{3N}\times[t_{i-1},t_{i})$. In general $dt$ does not represent a time-step (and the time-derivative is not approximated) and can often be taken much larger than usual time steps in standard computational solvers such as DMC. It is also possible to solve (2) by considering it as a boundary value problem, hence without approximating the time derivative. In this paper, as the objective is to propose links between PINN and N-VDMC algorithms, within PINN we propose a first order discretization of $\partial_{t}$, so that $dt$ actually represents the same time-step as the one used in DMC.
@@
    On $Q_{0}$ and for $f_{0}=\psi^{2}_{T}$ given, we search for a neural network $\widehat{f}:(\theta;R)\in\Theta\times\mathbb{R}^{3N}\rightarrow\mathbb{R}$ which i) optimized by minimizing the commonly called loss function
@@ 1.6 Inclusion of data
One of the key interests of neural networks is the use of data allowing to drive///supervise the learning the searched eigenstate(s). It basically consists in adding to the loss function, a data-fitting contribution allowing for a more accurate neural network optimization. As an example of supervised learning, let us consider a Schrödinger operator with parameter-dependent potentials $V(R,p)$ with $p\in\Pi$ and $\Pi$ a bounded subset of $\mathbb{R}^{\pi}$ with $\0\pi\in\mathbb{N}\backslash\{0\}$. Then we search for a neural network $\widehat{\phi}\,:\,(\theta;R,p)\in\Theta\times\mathbb{R}^{3N}\times\Pi\mapsto\widehat{\phi}(\theta;R,p)\in\mathbb{R}$ approximating a $p$-dependent groundstate $\phi_{G}\,:\,(R,p)\in\mathbb{R}^{3N}\times\Pi\mapsto\phi_{G}(R,p)\in\mathbb{R}$. Assuming that the groundstate was computed for $q$ parameters $p_{1},\cdots,p_{q}$, we add to the loss function the following contribution///constraint.
@@ 2.1 Preliminary remark
By abuse of notation (keeping the same notation for the training loss), and randomly selecting learning nodes $\{R^{\alpha}\}_{\alpha}$, we get
@@
Unlike PINN-DE, N-VDMC relies on the electron positions $\{R^{\alpha}_{i}\}_{\alpha}$
@@ 2.2 PINN as a fitting algorithm for DMC
The objective is here then to rewrite $\mathcal{L}^{\textrm{(PINN)}}_{i}$ using $\{f_{i}\}_{i}$, the N-VDMC solution. At iteration $i-1$, we assume given $\widehat{f}_{i-1}(R)$ the approximate solution to PINN-DE at time $t_{i}$. We select the smooth function $\Phi$, as follows
@@
In order to connect the PINN and N-VDMC algorithms, in the PINN loss we select the learning nodes $\{R^{\alpha}\}_{\alpha}$ over $Q_{i}=[t_{i},t_{i+1}]\times\mathbb{R}^{3N}$, as the electron positions from DMC; that is: $\{R^{\alpha}\}_{\alpha}=\{R_{i}^{\alpha}\}_{\alpha}$. For the sake of readiness, we use hereafter the following notation
$\left.\begin{array}[]{lll}\widehat{f}_{i-1}^{\alpha}=\widehat{f}_{i-1}(R_{i-1}^{\alpha}),&f_{i-1}^{\alpha}=f_{i-1}(R_{i-1}^{\alpha}),&f_{0;i}^{\alpha}=f_{0}(R_{i}^{\alpha}),\\ V_{i-1}^{\alpha}=\widehat{V}_{i-1}(R^{\alpha}_{i-1}),&\widehat{V}_{i-1}^{\alpha}\widehat{f}_{i-1}^{\alpha}=(\widehat{V}_{i-1}^{\alpha}\widehat{f}_{i-1}^{\alpha})_{i-1}^{\alpha},&V_{0;i-1}^{\alpha}=V_{0}(R^{\alpha}_{i-1})\\ \nabla_{R}\Phi_{i}^{\alpha}(\theta)=\nabla_{R}\Phi(\theta;R^{\alpha}_{i}),&\widehat{S}_{i-1}^{\alpha}=\widehat{S}_{i-1}(R^{\alpha}_{i-1}),&S_{0;i-1}^{\alpha}=S(R^{\alpha}_{i-1}),\\ \Phi_{i-1}^{\alpha}(\theta)=\Phi(\theta;R^{\alpha}_{i-1})\,.&&\end{array}\right.$
@@
We can actually refine this result. For $i=1$, as the initial data in PINN and DMC are taken identical then $\epsilon^{\alpha}_{0}$ and $\delta_{0}^{\alpha}$ are null, so that $\widehat{S}^{\alpha}_{0}=S^{\alpha}_{0;0}$, $\widehat{V}^{\alpha}_{0}=V^{\alpha}_{0;0}$ and as a consequence $\varepsilon_{0}^{\alpha}(\theta)=O(dt^{2})$. Hence, assuming that the loss can be as small as wanted (which is in general, a very strong assumption), we can deduce by induction that $\varepsilon_{i}^{\alpha}(\theta)=O(dt^{2})$. Thus, PINN can be seen as a data-fitting algorithm N-VDMC from data $\{f^{\alpha}_{i}\}$ up to a (small) correction $\varepsilon_{i-1}^{\alpha}$. More specifically, we can state the following proposition.
@@ 3 Optimization algorithm for Modified N-VDMC and PINN-DE
In this section we show that both the N-VDMC and PINN-DE algorithms lead at first order, to comparable optimization algorithm; hence lead to the comparable optimized neural network approximating the groundstate to the Schrödinger operator. More specifically, we show that the PINN method i) applied to the Imaginary Time Method (ITM) referred here as PINN-ITM, and ii) applied to the diffusion equation (2) PINN-DE, lead to a gradient descent algorithm similar to the one of a modified N-VDMC method. In the previous section, the N-VDMC algorithm (1.4) is initialized once for all at $t=t_{0}$ thanks to a N-VMC algorithm (Ferminet///Deep-Ritz). Hereafter, on each time interval $[t_{i},t_{i+1}]$, the so-called modified N-VDMC is reinitialized (at $t=t_{i}$) from a N-VMC algorithm, as long as $\widehat{\phi}_{i-1}$ is smooth and strictly positive. Beyond, this assumption we refer to Section 2. Moreover within the PINN algorithms, the learning nodes are taken as the electron positions from DMC.
@@
$\widehat{f}(\theta;\cdot)=\widehat{f}_{i-1}+dtK(\widehat{f}_{i-1}(\theta;\cdot))\,,$
@@
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\cfrac{dt}{2}\nabla_{R}(\widehat{\psi}\widehat{\phi}_{i-1})-dt\big{(}\nabla_{R}(\nabla_{R}|\log\widehat{\phi}_{i-1}|\widehat{\psi}\widehat{\phi}_{i-1})\big{)}+dtE_{T}\widehat{\psi}\widehat{\phi}_{i-1}-dt\widehat{\psi}H\widehat{\phi}_{i-1}\,.\end{array}\right.$
Some basic computations lead to
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\widehat{\psi}\cfrac{dt}{2}\nabla^{2}_{R}\widehat{\phi}_{i-1}+\cfrac{dt}{2}\widehat{\phi}_{i-1}\nabla^{2}_{R}\widehat{\psi}+dt\nabla_{R}\widehat{\psi}\nabla_{R}\widehat{\phi}_{i-1}\\ &&-dt\big{(}\nabla_{R}(\nabla_{R}\widehat{\psi}\widehat{\phi}_{i-1})-E_{T}\widehat{\psi}\widehat{\phi}_{i-1}+\widehat{\psi}H\widehat{\phi}_{i-1}\big{)}\,,\end{array}\right.$
@@
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\cfrac{dt}{2}\widehat{\phi}_{i-1}\nabla^{2}_{R}\widehat{\psi}-(dt-E_{T})\widehat{\psi}\widehat{\phi}_{i-1}\\ &=&\widehat{\phi}_{i-1}\big{(}\widehat{\phi}_{i-1}-dtH\widehat{\psi}\big{)}\,.\end{array}\right.$ (49)

### v2

@@ Abstract
In this note, we establish some connections between standard (data-driven) neural network-based solvers for PDE and eigenvalue problems developed on one side in the applied mathematics and engineering communities (e.g. Deep-Ritz and Physics Informed Neural Networks (PINN)), and on the other side in quantum chemistry (e.g. Variational Monte Carlo algorithms, Ferminet ([9]) or Paulinet ([15]) following the pioneer work of Carleo et. al [7, 8]. In particular, we re-formulate a PINN algorithm as a fitting problem with data corresponding to the solution to a standard Diffusion Monte Carlo algorithm initialized thanks to neural network-based Variational Monte Carlo. Connections at the level of the optimization algorithms are also established.
@@ 1 Introduction
In less than 10 years (since 2017), Physics Informed Neural Networks (PINN) has emerged as a major player in the field of computational methods for solving direct and inverse problems involving (high-dimensional) parameterized partial differential equations (PDEs) [23, 26, 19, 1]. The principle is to seek solutions to PDEs in the form of data-driven neural networks, optimized by minimizing PDE residuals. These techniques, along with others such as polynomial regression, Gaussian process, reduced order modeling, or Deep Ritz [25], have become highly popular in several engineering fields (e.g. mechanical engineering, fluid dynamics) as they allow to a certain extent, the tackling of problems that are inaccessible with traditional PDE solvers.
@@
Independently of the development of PINN-like methods by the mathematics, engineering, computer science, and statistics communities, significant practical progress has been achieved in quantum chemistry and quantum physics in the development of neural network-based Variational Monte Carlo (VMC) and Diffusion Monte Carlo (DMC) methods. This has enabled the calculation of the electronic structure of molecules larger than those usually considered with traditional computational methods. This approach corresponds mathematically to the approximation of the lowest eigenvalues (energies) of high-dimensional Schrödinger operators. Following the construction of Neural Quantum States (NQS) [7, 8], several corresponding libraries have been developed, such as Ferminet by DeepMind///Google [9, 22] or Paulinet ([15]).
To our knowledge, the two above formalisms have never been unified or even rigorously connected, despite their strong similarities. In this manuscript, we establish some connections potentially relevant for the development of a mathematical formalism unifying these modern techniques. More specifically, we will reformulate the so-called Neural-Variation Diffusion Monte Carlo (N-VDMC), corresponding roughly to a standard DMC method where the trial wavefunction is computed through a VMC method (corresponding to a Deep-Ritz algorithm [25]) as proposed for instance in Ferminet. Let us also mention [14], where a neural network based variational solving for many-body Schroedinger Hamiltonian is proposed. We will also show that at the level of the optimization algorithm (gradient descent-like algorithm) a N-VDMC is indeed “similar” to a PINN method applied to a diffusion equation (PINN-DE) derived from the time-dependent Schrödinger equation.
@@
In this manuscript, we do not discuss the (very complex) structure of the neural networks usually used in quantum chemistry which follows standard strategies based on one-electron orbitals and Slater’s determinants [24], and refer to [22, 7, 8] for details. However, the properties and connections presented in this manuscript are largely independent of the structure of the neural networks; the latter will be treated here as a parameterized functions.
@@ 1.2 Basics on neural networks
where $\sigma$ is a so-called activation function. We refer to [16, 2] for details on neural networks. The use of neural network-based algorithms for PDE and eigenvalue problems largely rely on approximability theorems such as the celebrated Cybenko’s Universal Approximation theorem [11] and more recently [12, 13]. Let us repeat that here, we do not discuss the very ellaborated structure of neural networks used for approximating wavefunctions in high-dimensional electronic structure problems, but refer to [22, 7, 8].
@@ 1.3 Deep-Ritz /// Ferminet
Following the celebrated works of Carleo et al. [7, 8] on neural quantum states, Ferminet [9, 22] has become a standard electronic structure calculation library, which is based on the minimization of the Rayleigh quotient
@@
hence defining an approximate groundstate. This approach can be seen as a Deep-Ritz algorithm [25]; corresponding to a class of variational neural network-based solvers for PDE and eigenvalue problems. In the following, the use of neural network as ansatz for minimizing the Rayleigh quotient will be referred as Neural-Variational Monte Carlo (N-VMC).
@@ 1.4 Diffusion Monte Carlo and Neural Diffusion Monte Carlo
leading to $E_{G}=\langle H\phi_{G},\psi_{T}\rangle/\langle\phi_{G},\psi_{T}\rangle$, as long as $\psi_{T}$ has an overlap with $\phi_{G}$. From a practical point of view the above approach is discretized using a sequence of time-dependent random walks $\{R^{\alpha}(t)\}_{\alpha}$ from $t$ to $\mathbb{R}^{3N}$, and allowing for the construction of a function $\widetilde{f}$ approximating $f$ in the form (see [27])
@@
A sequence of discrete times $t_{0}=0<t_{1}<\cdots<t_{i}<t_{i+1}<\cdots$ and $dt=t_{i+1}-t_{i}$ for any $i\in\mathbb{N}$ is introduced. For $\{R_{0}^{\alpha},t_{0}\}_{\alpha}$ given, we define $\widetilde{f}_{i}^{\alpha}=\widetilde{f}(R_{i}^{\alpha},t_{i})$, such that
@@
Finally, for $i$ large enough and denoting by $\langle\cdot\rangle_{\alpha}$ the average over the walkers, we get the mixed estimator at any time $t_{i}$, $E^{i}_{\textrm{mix}}=\langle\widehat{E}_{i-1}(R_{i}^{\alpha})\widetilde{f}_{i}^{\alpha}\rangle_{\alpha}/\langle\widetilde{f}_{i}^{\alpha}\rangle_{\alpha}$.
@@ 1.5 Basics on PINN
Physics Informed Neural Networks (PINN) [23, 26, 19, 1] has become a standard methodology in engineering and applied mathematics for the data-driven computation of partial differential equations. A PINN-like solution to a differential system ($\mathcal{S}$) is a (smooth and usually data-driven) neural network optimized thanks to the minimization of the norm of the PDE residual of ($\mathcal{S}$) evaluated at some (randomly chosen) learning nodes. This approach has recently gained a huge interest in scientific machine learning thanks to their simplicity, their nice mathematical and computational properties coming in particular from automatic differentiation and efficient stochastic gradient descent methods [5].
We now discuss the principle of PINN algorithms for solving (2). In order to avoid a null solution (which can occur for large times), it is preferable to decompose the time-domain as follows: $\mathbb{R}_{+}=\cup_{i=1}^{\infty}[t_{i-1},t_{i})$, with $t_{0}=0<t_{1}<\cdots<t_{i}<t_{i+1}<\cdots$ and we denote $dt=t_{i+1}-t_{i}$ and $Q_{i}=\mathbb{R}^{3N}\times[t_{i-1},t_{i})$. In general $dt$ does not represent a time-step (and the time-derivative is not approximated) and can often be taken much larger than usual time steps in standard computational solvers such as DMC. It is also possible to solve (2) by considering it as a boundary value problem, hence without approximating the time derivative. In this paper, as the objective is to propose links to N-VDMC algorithms, within PINN we propose a first order discretization of $\partial_{t}$, so that $dt$ actually represents the same time-step as the one used in DMC.
@@
    On $Q_{0}$ and for $f_{0}=\psi^{2}_{T}$ given, we search for a neural network $\widehat{f}:(\theta;R)\in\Theta\times\mathbb{R}^{3N}\rightarrow\mathbb{R}$ which i) optimized by minimizing the commonly called loss (objective) function
@@ 1.6 Inclusion of data
One of the key interests of neural networks is the use of data allowing to drive///supervise the learning of the searched eigenstate(s). It basically consists in adding to the loss function, a data-fitting contribution allowing for a more accurate neural network optimization. As an example of supervised learning, let us consider a Schrödinger operator with parameter-dependent potentials $V(R,p)$ with $p\in\Pi$ and $\Pi$ a bounded subset of $\mathbb{R}^{\pi}$ with $\0\pi\in\mathbb{N}\backslash\{0\}$. Then we search for a neural network $\widehat{\phi}\,:\,(\theta;R,p)\in\Theta\times\mathbb{R}^{3N}\times\Pi\mapsto\widehat{\phi}(\theta;R,p)\in\mathbb{R}$ approximating a $p$-dependent groundstate $\phi_{G}\,:\,(R,p)\in\mathbb{R}^{3N}\times\Pi\mapsto\phi_{G}(R,p)\in\mathbb{R}$. Assuming that the groundstate was computed for $q$ parameters $p_{1},\cdots,p_{q}$, we add to the loss function the following contribution///constraint.
@@ 2.1 Preliminary remark
By abuse of notation (keeping the same notation for the training loss), and for randomly selecting learning nodes $\{R^{\alpha}\}_{\alpha}$, we get
@@
Unlike PINN-DE, N-VDMC relies on the electron positions $\{R^{\alpha}_{i}\}_{\alpha}$ as follows
@@ 2.2 PINN as a fitting algorithm for DMC
The objective is here then to rewrite $\mathcal{L}^{\textrm{(PINN)}}_{i}$ using $\{f_{i}\}_{i}$, the N-VDMC solution. At iteration $i-1$, we assume given $\widehat{f}_{i-1}(R)$ the approximate solution to PINN-DE at time $t_{i}$. We construct the smooth function $\Phi$, as follows
@@
In order to connect the PINN and N-VDMC algorithms, in PINN’s loss function, we select the learning nodes $\{R^{\alpha}\}_{\alpha}$ over $Q_{i}=[t_{i},t_{i+1}]\times\mathbb{R}^{3N}$, as the electron positions from DMC; that is: $\{R^{\alpha}\}_{\alpha}=\{R_{i}^{\alpha}\}_{\alpha}$. For the sake of readiness, we use hereafter the following notation
$\left.\begin{array}[]{lll}\widehat{f}_{i-1}^{\alpha}=\widehat{f}_{i-1}(R_{i-1}^{\alpha}),&f_{i-1}^{\alpha}=f_{i-1}(R_{i-1}^{\alpha}),&f_{0;i}^{\alpha}=f_{0}(R_{i}^{\alpha}),\\ V_{i-1}^{\alpha}=\widehat{V}_{i-1}(R^{\alpha}_{i-1}),&\widehat{V}_{i-1}^{\alpha}\widehat{f}_{i-1}^{\alpha}=(\widehat{V}_{i-1}^{\alpha}\widehat{f}_{i-1}^{\alpha})_{i-1}^{\alpha},&V_{0;i-1}^{\alpha}=V_{0}(R^{\alpha}_{i-1})\\ \nabla_{R}\Phi_{i}^{\alpha}(\theta)=\nabla_{R}\Phi(\theta;R^{\alpha}_{i}),&\widehat{S}_{i-1}^{\alpha}=\widehat{S}_{i-1}(R^{\alpha}_{i-1}),&S_{0;i-1}^{\alpha}=S_{0}(R^{\alpha}_{i-1}),\\ \Phi_{i-1}^{\alpha}(\theta)=\Phi(\theta;R^{\alpha}_{i-1})\,.&&\end{array}\right.$
@@
We can actually refine this result. For $i=1$, as the initial data in PINN and DMC are taken identical then $\epsilon^{\alpha}_{0}$ and $\delta_{0}^{\alpha}$ are null, so that $\widehat{S}^{\alpha}_{0}=S^{\alpha}_{0;0}$, $\widehat{V}^{\alpha}_{0}=V^{\alpha}_{0;0}$ and as a consequence $\varepsilon_{0}^{\alpha}(\theta)=O(dt^{2})$. Hence, assuming that the loss can be as small as wanted (which is however in general, a very strong assumption), we can deduce by induction that $\varepsilon_{i}^{\alpha}(\theta)=O(dt^{2})$. Thus, PINN can be seen as a data-fitting algorithm N-VDMC from data $\{f^{\alpha}_{i}\}$ up to a (small) correction $\varepsilon_{i-1}^{\alpha}$. More specifically, we can state the following proposition.
@@ 3 Optimization algorithm for Modified N-VDMC and PINN-DE
In this section we show that both the N-VDMC and PINN-DE algorithms lead at first order, to comparable optimization algorithms; hence lead to comparable optimized neural networks approximating the groundstate to the Schrödinger operator. More specifically, we show that the PINN method i) applied to the Imaginary Time Method (ITM) referred here as PINN-ITM, and ii) applied to the diffusion equation (2) PINN-DE, lead to a gradient descent algorithm similar to the one of a modified N-VDMC method. In the previous section, the N-VDMC algorithm (1.4) is initialized once for all at $t=t_{0}$ thanks to a N-VMC algorithm (Ferminet///Deep-Ritz). Hereafter, on each time interval $[t_{i},t_{i+1}]$, the so-called modified N-VDMC is reinitialized (at $t=t_{i}$) from a N-VMC algorithm, as long as $\widehat{\phi}_{i-1}$ is smooth and strictly positive; we refer to Section 2. Moreover within the PINN algorithms, the learning nodes are taken as the electron positions from DMC. For simplicity, in the following we also set $E_{T}=0$.
@@
$\widehat{f}(\theta;\cdot)=\widehat{f}_{i-1}-dtK(\widehat{f}(\theta;\cdot))\,,$
@@
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\cfrac{dt}{2}\nabla^{2}_{R}(\widehat{\psi}\widehat{\phi}_{i-1})-dt\nabla_{R}\big{(}\widehat{\psi}\widehat{\phi}_{i-1}\nabla_{R}|\log\widehat{\phi}_{i-1}|\big{)}-dt\widehat{\psi}H\widehat{\phi}_{i-1}\,.\end{array}\right.$
Then we get
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\widehat{\psi}\cfrac{dt}{2}\nabla^{2}_{R}\widehat{\phi}_{i-1}+\cfrac{dt}{2}\widehat{\phi}_{i-1}\nabla^{2}_{R}\widehat{\psi}+dt\nabla_{R}\widehat{\psi}\nabla_{R}\widehat{\phi}_{i-1}-dt\nabla_{R}(\widehat{\psi}\nabla_{R}\widehat{\phi}_{i-1})-dt\widehat{\psi}H\widehat{\phi}_{i-1}\,,\end{array}\right.$
@@
$\left.\begin{array}[]{lcl}\widehat{\psi}\widehat{\phi}_{i-1}&=&\widehat{\phi}^{2}_{i-1}+\cfrac{dt}{2}\widehat{\phi}_{i-1}\nabla^{2}_{R}\widehat{\psi}-\cfrac{dt}{2}\widehat{\psi}\nabla^{2}_{R}\widehat{\phi}_{i-1}-dt\widehat{\phi}_{i-1}V_{p}(R)\widehat{\psi}+\cfrac{dt}{2}\widehat{\psi}\nabla^{2}_{R}\widehat{\phi}_{i-1}\\ &=&\widehat{\phi}_{i-1}\big{(}\widehat{\phi}_{i-1}-dtH\widehat{\psi}\big{)}\,.\end{array}\right.$ (49)